### v1

@@ 3 PolyDG formulation
$\left\{\begin{aligned} &\mathscr{K}_{h}\mathbf{U}^{(k+1)}=\left(\chi_{m}C_{m}\mathbf{M}_{h}-\frac{\Delta t}{2}\mathbf{A}_{h}\right)\mathbf{U}^{(k)}-\mathrm{\chi_{m}}\Delta t\mathbf{I}^{(k+1)}+\Delta t\mathbf{F}^{(k+1)},\\ &\mathbf{Y}^{(k+1)}=\mathbf{Y}^{(k)}-\Delta t\mathbf{G}^{(k)},\\ &\mathbf{U}^{0}=\mathbf{U}_{0},\;\mathbf{Y}^{0}=\mathbf{Y}_{0}.\end{aligned}\right$. (4)
The ionic model can be efficiently solved (since $\mathbf{M}_{h}$ is block diagonal). On the other hand, the monodomain equation involving $\mathscr{K}_{h}$ has to be solved at each time step by a suitable Krylov-type solver (e.g., conjugate gradient). In the next section, we will present a non-overlapping Schwarz preconditioner to accelerate its convergence.

### v2

@@ 3 PolyDG formulation
$\left\{\begin{aligned} &\mathscr{K}_{h}\mathbf{U}^{(k+1)}=\left(\chi_{m}C_{m}\mathbf{M}_{h}-\frac{\Delta t}{2}\mathbf{A}_{h}\right)\mathbf{U}^{(k)}-\mathrm{\chi_{m}}\Delta t\mathbf{I}^{(k+1)}+\Delta t\mathbf{F}^{(k+1)},\\ &\mathbf{M}_{h}\mathbf{Y}_{l}^{(k+1)}=\mathbf{M}_{h}\mathbf{Y}_{l}^{(k)}-\Delta t\mathbf{G}_{l}^{(k)},&\forall l=1,\cdots,n\;,\\ &\mathbf{U}^{0}=\mathbf{U}_{0},\;\mathbf{Y}^{0}=\mathbf{Y}_{0}.\end{aligned}\right$. (4)
where $n$ is the dimension of the Barreto-Cressman ionic model that can be efficiently solved (since $\mathbf{M}_{h}$ is block diagonal). On the other hand, the monodomain equation involving $\mathscr{K}_{h}$ has to be solved at each time step by a suitable Krylov-type solver (e.g., conjugate gradient). In the next section, we will present a non-overlapping Schwarz preconditioner to accelerate its convergence.